Physics problem:

What is the minimum horizontal force that will cause a 5.00-kg box to being to slide on a horizontal surface when the coefficient of static friction is 0.670?
Answer #1

minimum force required is the frictional force

frictional force = mu_s*N = mu_s*m*g

mu_s = 0.67

m = 5 kg

g = 9.8 m/s^2

frictional force = 0.67*5*9.8 =32.83 N....is the required answer.

and the force greater than 32.83 N causes motion of the body